OpenStudy (anonymous):

(a) −3 and 6 are the first two terms of an arithmetic progression. Write down the constant difference, calculate the 23rd term, and find the sum of the first 23 terms. (b) Which term an of this A.P. is equal to 276?

OpenStudy (anonymous):

@ghass1978 @jim_thompson5910

OpenStudy (anonymous):

Hi the difference id clearly their difference, i.e. 9

OpenStudy (anonymous):

\[a_{23} = a1 + (23-1)* d = -3 + 22 * 9\]

OpenStudy (anonymous):

S_23 = (a1 + a23 )*23/2 = 91*23

OpenStudy (anonymous):

Do U understand ?

OpenStudy (anonymous):

Ummm, not really... I barely understand the question...

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

@eseidl

OpenStudy (anonymous):

@Mikael